# Cosmic Cartography II: completing galaxy catalogs for gravitational-wave cosmology

### v1

@@ 4.4 Impact of the absolute magnitude threshold
Our inferred absolute magnitude distribution is faithful only up to a given threshold, determined by the faintest near-by galaxies. [^10] Note that this assumes that the absolute magnitude distribution only weakly varies with redshift. Hence, when we infer the galaxy Poisson rate in each four-dimensional redshift, sky position and absolute magnitude bin, we set all counts above an absolute magnitude threshold, denoted as $M_{\rm thr}$, to zero. Ideally, this threshold coincides with the absolute magnitude of the faintest galaxy that has a non-negligible GW host probability. [^11] In practice, this is difficult, as the galaxy GW host probability is not understood. If the probability of hosting a GW event is negligible for all galaxies above some $M_{\rm thr}$ (as is expected from scaling relations relating star formation rates to the overall galaxy luminosity [79]), we can safely ignore all galaxies with $M\geq M_{\rm thr}$. This section varies this threshold $M_{\rm thr}$ and studies its impact on the posterior of the galaxy Poisson rate.
@@ 5 Conclusions
Gravitational waves from CBCs are a promising avenue to measure the cosmic expansion history independently of traditional methods that solely use electromagnetic information. To constrain the cosmological parameters with GWs, the main bottleneck is obtaining accurate redshift information of the CBC. The dark siren method addresses this. It leverages galaxy catalogs into a redshift measurement by “overlaying” the GW sky position with the catalog. As such, the galaxy catalog provides a redshift-sky position prior for the GW events. However, current methods that combine the GW data with galaxy catalogs only slightly improve the $H_{0}$ precision when compared to using GW data alone. This is mostly due to the limitations of the employed catalogs, which only contain galaxies up to $z\sim 0.3$ – galaxies above that redshift are too faint to be confidently observed. To address this, this work exploits the spatial correlations (through large-scale structure) – not used in the standard codes – to reconstruct the full galaxy density field. Our reconstruction improves the redshift–sky position prior for the GW signal, and through the more accurate redshift prior for GW events will lead to more informative constraints on the Hubble constant, something we plan to to demonstrate in the near future.
We have demonstrated a Bayesian approach to reconstruct a galaxy catalog from a magnitude limited survey. To this end, we have simultaneously reconstructed the galaxy magnitude distribution, the DM power spectrum, the DM density contrast, the DM-galaxy bias, the Hubble constant and selection effects. All the above ingredients are indispensable to connect the observed catalog to the full galaxy field. We have demonstrated that one can recover the original (true) catalog to a high degree, and that the inferred uncertainties in the galaxy rate are accurate. Through our agnostic Gaussian process magnitude distribution, our method does not require a priori information about the galaxy absolute magnitude distribution or its parametric form.
@@ A.1 Gaussian Magnitude Distribution
[…]
@@ Appendix B Interpolation from cartesian to spherical coordinates
To be able to use the existing infrastructure of the log-normal field from the previous project [65] based on [67, 68], this section summarizes how we connect the field in cartesian and spherical coordinates. For densities, the field value of the spherical element is then simply the average of the field values of the cartesian elements that fall inside that spherical element. Therefore, we require that at least one center of a cartesian volume element to fall inside each of the three-dimensional spherical volume elements. For a given spherical discretization this implies the minimum density of cartesian elements. However, since the spherical elements are densely packed at the origin, this would require a very fine cartesian grid, with many cartesian elements in the region far from the origin. Thus, we refine this method in the following way: centered at the origin, we define a cubic region with side length $L/2$ (for $L$ the length of the total cube) at which all cartesian elements are repeated twice along each dimension. See Fig. 20 that illustrates the denser three-dimensional grid in the central region of the simulated cube.

### v3

@@ 4.4 Impact of the absolute magnitude threshold
Our inferred absolute magnitude distribution is faithful only up to a given threshold, determined by the faintest near-by galaxies. [^9] Note that this assumes that the absolute magnitude distribution only weakly varies with redshift. Hence, when we infer the galaxy Poisson rate in each four-dimensional redshift, sky position and absolute magnitude bin, we set all counts above an absolute magnitude threshold, denoted as $M_{\rm thr}$, to zero. Ideally, this threshold coincides with the absolute magnitude of the faintest galaxy that has a non-negligible GW host probability. [^10] In practice, this is difficult, as the galaxy GW host probability is not understood. If the probability of hosting a GW event is negligible for all galaxies above some $M_{\rm thr}$ (as is expected from scaling relations relating star formation rates to the overall galaxy luminosity [Artale:2019tfl]), we can safely ignore all galaxies with $M\geq M_{\rm thr}$. This section varies this threshold $M_{\rm thr}$ and studies its impact on the posterior of the galaxy Poisson rate.
@@ 5 Conclusions
Gravitational waves from CBCs are a promising avenue to measure the cosmic expansion history independently of traditional methods that solely use electromagnetic information. To constrain the cosmological parameters with GWs, the main bottleneck is obtaining accurate redshift information of the CBC. The dark siren method addresses this. It leverages galaxy catalogs into a redshift measurement by “overlaying” the GW sky position with the catalog. As such, the galaxy catalog provides a redshift-sky position prior for the GW events. However, current methods that combine the GW data with galaxy catalogs only slightly improve the $H_{0}$ precision when compared to using GW data alone. This is mostly due to the limitations of the employed catalogs, which only contain galaxies up to $z\sim 0.3$ – galaxies above that redshift are too faint to be confidently observed. To address this, this work exploits the spatial correlations (through large-scale structure) – not used in the standard codes – to reconstruct the full galaxy density field. Our reconstruction improves the redshift–sky position prior for the GW signal, and through the more accurate redshift prior for GW events will lead to more informative constraints on the Hubble constant, something we plan to demonstrate in the near future.
We have demonstrated a Bayesian approach to reconstruct a galaxy catalog from a magnitude limited survey. To this end, we have simultaneously reconstructed the galaxy magnitude distribution, the DM power spectrum, the DM density contrast, the DM-galaxy bias, the Hubble constant and selection effects. All the above ingredients are indispensable to connect the observed catalog to the full galaxy field. We have demonstrated that one can recover the original (true) catalog to a high degree, and that the inferred uncertainties in the galaxy rate are accurate. Through our agnostic Gaussian process magnitude distribution, our method does not require a priori information about the galaxy absolute magnitude distribution or its parametric form. This is with the caveat that this distribution was assumed to be independent of the DM density contrast –the effect of high density clusters producing brighter galaxies was neglected in this approach.
@@ A.1 Gaussian Magnitude Distribution
and we assume a small regularization value $\text{cut-off}=10^{-6}$. In this expression:
@@ Appendix B Interpolation from cartesian to spherical coordinates
To be able to use the existing infrastructure of the log-normal field from the previous project [Leyde:2024tov] based on [Coles:1991if, Agrawal:2017khv], this section summarizes how we connect the field in cartesian and spherical coordinates. For densities, the field value of the spherical element is then simply the average of the field values of the cartesian elements that fall inside that spherical element. Therefore, we require that at least one center of a cartesian volume element to fall inside each of the three-dimensional spherical volume elements. For a given spherical discretization this implies the minimum density of cartesian elements. However, since the spherical elements are densely packed at the origin, this would require a very fine cartesian grid, with many cartesian elements in the region far from the origin. Thus, we refine this method in the following way: centered at the origin, we define a cubic region with side length $L/2$ (for $L$ the length of the total cube) at which all cartesian elements are repeated twice along each dimension. See Fig. 20 that illustrates the denser three-dimensional grid in the central region of the simulated cube.